46% of U.S. adults have very little confidence in newspapers. You randomly select 10 U.S. adults. Find the probability that t46​% of U.S. adults have very little confidence in newspapers. You randomly select 10 U.S. adults. Find the probability that the number of U.S. adults who have very little confidence in newspapers is ​ (a) exactly​ five, (b) at least​ six, and​ (c) less than four.
Answer #1

X ~ Binomial (n,p)

Where n = 10 , p = 0.46

P(X) = nCx px (1-p)n-x

a)

P(5) = 10C5 0.465 0.545

= 0.238

b)

P( X >= 6) = P (X = 6) + P (X = 7) + P (X = 8) + P (X = 9) + P (X = 10)

= 10C6 0.466 0.544 +10C7 0.467 0.543 +10C8 0.468 0.542 +10C9 0.469 0.541 +10C10 0.4610 0.540

= 0.283

c)

P( X < 4) = P( X <= 3)

= P( X = 0) + P( X = 1 ) + P(X = 2 ) + P (X = 3)

= 10C0 0.460 0.5410 +10C1 0.461 0.549 +10C2 0.462 0.548 +10C3 0.463 0.547

= 0.245